# Small but large: Single organic molecules as hybrid platforms for quantum technologies

### v1

@@ Abstract
Single organic molecules embedded in solid-state matrices exhibit remarkable optical properties, making them competitive candidates for single-photon sources and quantum nonlinear optical elements. However, the lack of long-lived internal states imposes significant constraints on their application in the broader field of quantum technologies. In this article, we reexamine the single-molecule host-guest system from first principles, elaborate on the rich internal states this system encompasses and put forward strategies to harness them for applications in quantum memory, spin-photon interface, spin register, and optomechanics. Further, we discuss the potential of leveraging the vast chemical space of molecules, and highlight the challenges and opportunities for molecular systems along these directions.
@@ I Introduction
Since the first detection of a single pentacene molecule in para-terphenyl crystal by Moerner et al. [1] and Orrit et al. [2], the favorable optical properties of polycyclic aromatic hydrocarbon (PAH) molecules have enabled many pioneering advancements in quantum optics [3, 4]. Their preeminent properties of single photon emission, such as the Fourier-limited linewidth, high quantum efficiency, and high photostability make them excellent sources of indistinguishable photons [5, 6, 7]. A single molecule can function as an optical nonlinear medium operating in the quantum regime, owing to the anharmonicity of its electronic excited states. This has been utilized to demonstrate a single-molecule optical transistor [8] and frequency mixing between two optical beams [9, 10]. Furthermore, the frequency of molecular photons can be tuned by the Stark effect [11, 12], exciting with a strong pump laser [13] or external strain [14, 15, 16]. The fine control over the emission frequency facilitated the interfacing of remote molecules through single photons [17, 18]. Moreover, PAH molecules are compatible with photonic structures such as dielectric [19, 20] and plasmonic antennas [21], optical waveguides [22, 23, 24, 25, 26] and on-chip resonators [27]. Recent demonstrations have shown the strong coupling of a single molecule to a Fabry-Pérot microcavity [18, 10]. Having frequency tunability and controllable doping concentration, near-field couplings between two closely spaced molecules via coherent dipole-dipole interactions have also been shown [28, 29, 30]. In addition, the narrow-linewidth optical transition makes single molecules ultra-sensitive probes for their local environment, e.g., charge fluctuations [31], and mechanical oscillations [32, 15].
@@
Spectroscopic detection of single PAH molecules relies on their insertion as impurities into a host material, i.e., forming a host-guest system through van der Waals interactions. With suitable choices of host-guest combinations, such a system can feature two favorable properties. First, the van der Waals interaction allows the electronic excitations of the guest to remain largely localized within its own molecular orbitals, with only weak perturbation from the host. In this regard, the guest’s ground state and lowest electronic excited state facilitate an optical transition, which could feature a nearly lifetime-limited linewidth at liquid helium temperatures, on par with atomic transitions in vacuum [3]. Second, this transition can be approximated as a two-level cycling transition, where population decays to other levels are recycled through interactions with the matrix, e.g., fast vibrational relaxation, allowing measurements using only a single laser beam. It is worth noting that the success of most past achievements in molecular quantum optics builds primarily upon the efficient isolation of such a two-level transition. Beyond this simplified two-level picture, the molecular host-guest system represents a hybrid quantum system, incorporating electronic, nuclear, and spin degrees of freedom [33]. The presence of this multitude of couplings has been considered detrimental, as they introduce incoherent leakage channels.
@@
In this article, we introduce single molecules embedded in a solid-state host as a hybrid quantum system, provide a detailed discussion of its various degrees of freedom, and outline the prospects of harnessing them for applications that have not been realized on molecular systems so far. The article is organized as follows. In section II, we examine the molecular host-guest system from first principles, revisit its wavefunction, energy landscape, and associated transitions. A particular focus will be given to the quantum nature of its vibrational and spin degrees of freedom. In section III, we discuss the potential of using vibrational and spin states for molecular quantum memory, optomechanics, spin-photon interfaces and chemically configurable spin registers. We will further expand the discussion towards exploring the vast chemical space of molecules for quantum applications. Finally, we discuss the theoretical and experimental challenges associated with these approaches in section IV and conclude with section V.
@@
Figure 1: Upper panel: Illustration of a molecular host-guest system, specifically shown here for anthracene crystal hosting a perylene molecule. Lower Panel: Schematics for the relevant states and their couplings in a molecular host-guest system. The guest’s spin-orbital states (top) interact with its own intramolecular vibrations (bottom left) and intermolecular vibrations of the host crystal (bottom right) via electron-vibron and electron-phonon couplings, respectively. Electronic states are further coupled to nuclear spin degrees of freedom via hyperfine interactions (middle). The vibrational states are coupled to each other via vibron-phonon couplings.
@@ II Emerging Insights into Molecular Host-Guest Systems
Molecules have emerged as engineerable building blocks for future quantum technologies, thanks to their rich internal structures and the vast diversity offered by the chemical space [34, 35, 3, 36]. Figure 1 illustrates the multiple degrees of freedom in the molecular host-guest system and their couplings. While there is also growing interest in exploiting these internal states [37, 38, 39, 40], earlier studies have predominantly focused on the electronic degree of freedom [3], leaving other intriguing aspects largely unexplored. To highlight the promise of molecular host-guest systems for future explorations in quantum science, we revisit these “overlooked” internal states, with a focus on their quantum mechanical character and interactions. Drawing inspiration from previous works [41], we perform a first-principle analysis of molecular host-guest systems to explore their quantum mechanical character in detail.
@@ II.1 Revisiting the Host-Guest Wavefunction
where $\mathbf{r}_{\text{g}}$,$\mathbf{s}_{\text{g}}$ ($\mathbf{r}_{\text{h}}$,$\mathbf{s}_{\text{h}}$) denote the spatial and spin coordinates of the guest (host), respectively. When studying low-lying excitations in the guest, the electronic states of the host are unaffected and can be traced out  [^1] Note that the host’s electronic degrees of freedom plays an important role in determining the geometry of the guest molecule in the solid state.. Hence, the system’s electronic wavefunction reduces to the spin-orbital state of the guest $\Psi_{\text{el,guest}}$ (see lower panel of Fig. 1).
@@
Moreover, decoupling of the host and guest molecular orbitals allows the decomposition of the nuclear wavefunction into those of the host and the guest. The composite nuclear wavefunction can be expressed in terms of he normal modes of the guest molecule and the host matrix with coordinates $\mathbf{Q}$ and $\mathbf{u}$ as [41, 42]
@@
where $\psi_{\text{nuc,guest}}(\mathbf{Q})$ and $\psi_{\text{nuc,host}}(\mathbf{u})$ denote the intramolecular guest and intermolecular host vibrational wavefunctions (see bottom part of Fig. 1). In the remainder of the text, we use the terms vibron for the quantized intramolecular normal modes of guest molecule and phonon for the quantized intermolecular normal modes of host matrix.
@@
with their order reflecting the hierarchy of energy scales associated with these states.
@@
where $j$ is the integer index of the spin-orbital state. In the following discussion of electronic and vibronic transitions, we will first omit the nuclear spin states, as they primarily manifest through hyperfine interactions with electron spins and are relevant only when probing the triplet manifold, such as in optically detected magnetic resonance experiments. We will revisit these states in the discussion of spin states.
Figure 2: (a) The energy levels of a single molecule embedded in a matrix. The state |$\ket{S_{0(1)};\bm{\nu}_{v},\bm{\nu}_{p}}$ belongs to the ground (excited) state manifold and has vibrational occupations in the intramolecular |$\ket{\bm{\nu}_{v}}=\Pi_{i}\ket{\nu_{v}^{(i)}}$ and intermolecular vibrational modes |$\ket{\bm{\nu}_{p}}=\Pi_{i}\ket{\nu_{p}^{(i)}}$. The indices $\nu_{v(p)}^{(i)}=0,1\ldots$ denote the vibrational quantum number of the $i^{\text{th}}$ vibron (phonon) mode. Similarly, |$\ket{T_{1};\bm{\nu}_{v},\bm{\nu}_{p}}$ represents the states in the triplet state manifold. (b) PAH molecules commonly used in quantum optics experiments and their range of approximated zero-phonon-line wavelengths [43].
@@ II.1.1 Ground and excited states
The ground state of the system |$\ket{S_{0};\bm{\nu}_{v}=\mathbf{0},\bm{\nu}_{p}=\mathbf{0}}$ is a singlet state, in which all $\pi$-orbitals are filled and all electron spins are paired. The excitation of one electron to a $\pi^{*}$ orbital leaves the molecule in the lowest singlet excited state |$\ket{S_{1};\bm{\nu}_{v}=\mathbf{0},\bm{\nu}_{p}=\mathbf{0}}$. Alternatively, if this excitation is accompanied by a spin-flip, the molecule reaches the lowest triplet excited state |$\ket{T_{1};\bm{\nu}_{v}=\mathbf{0},\bm{\nu}_{p}=\mathbf{0}}$. The energy of |$\ket{T_{1};\bm{\nu}_{v}=\mathbf{0},\bm{\nu}_{p}=\mathbf{0}}$ is lower than that of |$\ket{S_{1};\bm{\nu}_{v}=\mathbf{0},\bm{\nu}_{p}=\mathbf{0}}$ by the exchange energy. The transition energies of these two excited states to the ground state lie in the optical frequency domain and range from UV to near-infrared wavelengths, depending on the size and geometry of the molecule. The triplet state |$\ket{T_{1};\bm{\nu}_{v}=\mathbf{0},\bm{\nu}_{p}=\mathbf{0}}$ consists of three sublevels, resulting from the anisotropic magnetic dipole-dipole interactions of two unpaired electrons, which remain non-degenerate even in the absence of an external magnetic field [44]. This is called zero-field splitting (ZFS). For PAH molecules emitting in the UV to visible range, the ZFS can reach a few gigahertz [45]. Moreover, coupling of the electron spin to nuclear spins in the molecule and the matrix leads to hyperfine splittings, which lie in the radio-frequency range.
@@
In addition, each manifold includes vibron states |$\ket{S(T),\bm{\nu}_{v}\neq\mathbf{0},\bm{\nu}_{p}=\mathbf{0}}$ corresponding to librational and bond vibrations of the guest molecule, with frequencies that can reach up to $100$ THz. The coupling between electronic and intramolecular vibrational states (electron-vibron coupling) gives rise to the well-known Franck-Condon physics, where the displacement of the excited state potential energy surface along a normal-mode coordinate enables transitions between vibronic states in different manifolds. The transition probabilities, i.e., Franck-Condon factors, are related to the Huang-Ryhs factor of each normal mode [46]. For certain PAH molecules, the simple Franck-Condon picture can be further complicated by the Herzberg-Teller effect, where couplings between the electronic and nuclear degrees of freedom are not parametric [47, 48]. Furthermore, non-adiabatic couplings between the electronic and nuclear degrees of freedom result in non-radiative decay from the excited state manifold via electron-vibron couplings [49, 50]. This could lead to non-unity quantum efficiencies of photon emission [51, 20].
@@
Besides, phonon states |$\ket{S(T),\bm{\nu}_{v}=\mathbf{0},\bm{\nu}_{p}\neq\mathbf{0}}$ typically fall within the $0-5$ THz range and lead to the deformation of molecular orbitals. The resulting electron-phonon couplings yield various optical transitions between different manifolds involving vibrons and phonons, similar to the Franck-Condon physics discussed earlier. Depending on the frequency range, these vibrational states exhibit lifetimes in the range of milliseconds to picoseconds due to different vibrational relaxation mechanisms [52].
@@ II.1.2 Zero-phonon-line transition
In summary, a single molecule embedded in a solid-state host exhibits a rich energy structure supporting a multitude of well-defined transitions ranging from radio to optical frequencies. Next, we delve into the details of these molecular degrees of freedom and discuss their quantum mechanical character.
@@ II.2 Vibrational States
The optomechanical character of molecules was first discovered by Raman [55], as the optical excitation of molecules yields light scattering to new frequencies determined by the intramolecular vibrational energies. In the context of molecular quantum optics, this optomechanical character has traditionally been regarded as a source of decoherence due to fast vibrational relaxation processes. Yet, if vibrational states can be engineered to have long lifetimes, they could be exploited as a resource in quantum technologies.
Figure 3: (a) Phonon density of states (DOS) of naphthalene, and anthracene crystal (top) and intramolecular vibrations of a perylene molecule (bottom), shown by red, black and green lines, respectively. The ab initio calculations are done using Vienna Ab initio Simulation Package [56, 57, 58, 59] with Perdew-Burke-Ernzerhof functional [60] and Tkatchenko-Scheffler dispersion correction [61]. We used phonopy to calculate the phonon DOS for a $2\times 4\times 2$ supercell [62] (b, c) Schematics for the acoustic and optical phonon of an anthracene crystal and pseudolocal mode of the host-guest system. The crystal axes are denoted by $a$ and $b$, respectively. (d) Vibrational relaxation mechanisms for an intramolecular vibration of a guest molecule in a host environment. $\omega_{\nu_{v}}$ and $\omega_{\text{max}}$ denote the intramolecular vibrational frequency of the state |$\ket{S_{0};[0,\ldots,0,1,0\dots,0],\bm{\nu}_{p}=\bm{0}}$ and maximal phonon frequency in the host matrix, respectively.

### v2

@@ Abstract
Single organic molecules embedded in solid-state matrices exhibit remarkable optical properties, making them competitive candidates for single-photon sources and quantum nonlinear optical elements. However, the lack of long-lived internal states imposes significant constraints on their application in the broader field of quantum technologies. In this article, we reexamine the single-molecule host-guest system from first principles, elaborate on the rich internal states this system encompasses and put forward strategies to harness them for applications in quantum memory, spin-photon interface, spin register, and optomechanics. Further, we discuss the potential of leveraging the vast chemical space of molecules and highlight the challenges and opportunities for molecular systems along these directions.
@@ I Introduction
Molecular systems have emerged as promising platforms for quantum technologies due to their unique combination of structural tunability, rich internal degrees of freedom, and compatibility with solid-state integration. A notable example of their potential is the detection of a single pentacene molecule in a para-terphenyl crystal by Moerner et al. [1] and Orrit et al. [2], which marked the beginning of significant advancements in quantum optics driven by polycyclic aromatic hydrocarbon (PAH) molecules [3, 4]. Over time, the field of solid-state molecular quantum technologies has expanded to include a range of spin-based platforms, such as organometallic systems [5], rare-earth molecular complexes [6] and organic radical species [7], driven by the versatility offered through chemical synthesis.
@@
Among these systems, single PAH molecules are distinguished by their preeminent properties, including single-photon emission, close to Fourier-limited linewidth, high quantum efficiency, and excellent photostability, which make them excellent sources of indistinguishable photons [8, 9, 10]. These characteristics also enable single molecules to function as optical nonlinear media operating in the quantum regime, owing to the anharmonicity of their electronic excited states. This has been utilized to demonstrate a single-molecule optical transistor [11] and frequency mixing between two optical beams [12, 13]. Furthermore, the frequency of molecular photons can be tuned by the Stark effect [14, 15, 16], exciting with a strong pump laser [17] or external strain [18, 19, 20]. The fine control over the emission frequency facilitated the interfacing of remote molecules through single photons [21, 22]. Moreover, PAH molecules are compatible with photonic structures such as dielectric [23, 24] and plasmonic antennas [25], optical waveguides [26, 27, 28, 29, 30] and on-chip resonators [31]. Recent experiments have demonstrated the strong coupling of a single molecule to a Fabry-Pérot microcavity [22, 13]. Having frequency tunability and controllable doping concentration, near-field couplings between two closely spaced molecules via coherent dipole-dipole interactions have also been shown [32, 33, 34]. In addition, the narrow-linewidth optical transition makes single molecules ultra-sensitive probes for their local environment, e.g., charge fluctuations [35], and mechanical oscillations [36, 19].
@@
Spectroscopic detection of single PAH molecules relies on their insertion as impurities into a host material, i.e., forming a host-guest system through van der Waals interactions. With suitable choices of host-guest combinations, such a system can feature two favorable properties. First, the van der Waals interaction allows the electronic excitations of the guest to remain largely localized within its own molecular orbitals, with only weak perturbation from the host. In this regard, the guest’s ground state and lowest electronic excited state facilitate an optical transition, which could feature a nearly lifetime-limited linewidth at liquid helium temperatures, on par with atomic transitions in vacuum. Second, this transition can be approximated as a two-level cycling transition, where population decays to other levels are recycled through interactions with the matrix, e.g., fast vibrational relaxation, allowing measurements using only a single laser beam. It is worth noting that the success of most past achievements in molecular quantum optics builds primarily upon the efficient isolation of such a two-level transition. Beyond this simplified two-level picture, the molecular host-guest system represents a hybrid quantum system, incorporating electronic, nuclear, and spin degrees of freedom. The presence of this multitude of couplings has been considered detrimental, as they introduce incoherent leakage channels.
@@
In this article, we introduce single molecules embedded in a solid-state host as a hybrid quantum system, provide a detailed discussion of their various degrees of freedom, and outline the prospects of harnessing them for applications that have not been realized on molecules so far. The article is organized as follows. In section II, we examine the molecular host-guest system from first principles, revisit its wavefunction, energy landscape, and associated transitions. A particular focus will be given to the quantum nature of its vibrational and spin degrees of freedom. In section III, we discuss the potential of using vibrational and spin states for molecular quantum memory, optomechanics, spin-photon interfaces and chemically configurable spin registers. We will further expand the discussion towards exploring the vast chemical space of molecules for quantum applications. Finally, we discuss the theoretical and experimental challenges associated with these approaches in section IV and conclude with section V.
@@
Figure 1: Upper panel: Illustration of a molecular host-guest system, specifically shown here for anthracene crystal hosting a perylene molecule. Lower Panel: Schematics for the relevant states and their couplings. The guest’s spin-orbital states (top) interact with its own intramolecular vibrations (lower left) and intermolecular vibrations of the host crystal (lower right) via electron-vibron and electron-phonon couplings, respectively. Electronic states are further coupled to nuclear spin degrees of freedom via hyperfine interactions (middle). The vibrational states are coupled to each other via vibron-phonon couplings. The energy hierarchy of these states is shown at the bottom of the lower panel, where $E_{\text{electronic}}\sim$ 1 eV - 3 eV, $E_{\text{vibron}}\sim$ 20 meV - 0.5 eV, $E_{\text{phonon}}\sim$ 0.3 neV - 20 meV, $E_{\text{nuclear spin}}\sim$ $0.4$ neV - $400$ neV. The quoted energy scale for nuclear spins refers to their range of hyperfine coupling strength.
@@ II Emerging Insights into Molecular Host-Guest Systems
Molecules have emerged as engineerable building blocks for future quantum technologies, thanks to their rich internal structures and the vast diversity offered by the chemical space [37, 38, 3, 39]. Figure 1 illustrates the multiple degrees of freedom in the molecular host-guest system and their couplings. While there is also growing interest in exploiting these internal states [40, 41, 42], earlier studies have predominantly focused on the electronic degree of freedom [3], leaving other intriguing aspects largely unexplored. To highlight the promise of molecular host-guest systems for future explorations in quantum science, we revisit these “overlooked” internal states. Drawing inspiration from previous works [43], we perform a first-principle analysis of molecular host-guest systems to explore their quantum mechanical characters and interactions in detail.
@@ II.1 Revisiting the Host-Guest Wavefunction
where $\mathbf{r}_{\text{g}}$, $\mathbf{s}_{\text{g}}$ ($\mathbf{r}_{\text{h}}$, $\mathbf{s}_{\text{h}}$) denote the spatial and spin coordinates of the guest (host), respectively. When studying low-lying excitations in the guest, the electronic states of the host are unaffected and can be traced out  [^1] The host’s electronic degrees of freedom dress that of the guest molecule, and play an important role in determining the geometry of the guest molecule in the solid state.. Hence, the system’s electronic wavefunction reduces to the spin-orbital state of the guest $\Psi_{\text{el,guest}}$, as shown by the top node of the triangular diagram in Fig. 1.
@@
Moreover, decoupling of the host and guest molecular orbitals allows the decomposition of the nuclear wavefunction into those of the host and the guest. The vibrational part of the nuclear wavefunction can be expressed in terms of the normal modes of the guest molecule and the host matrix with coordinates $\mathbf{Q}$ and $\mathbf{u}$ as [43, 44]
@@
where $\psi_{\text{nuc,guest}}(\mathbf{Q})$ and $\psi_{\text{nuc,host}}(\mathbf{u})$ denote the intramolecular guest and intermolecular host vibrational wavefunctions, respectively (see the bottom nodes of the triangular diagram in Fig. 1.) We note that the separation between host and guest vibrational degrees of freedom breaks down for low-frequency guest modes, which appear in large PAH molecules. These modes are discussed in detail in Section II.2.2. In the remainder of the text, we use the terms vibron for the quantized intramolecular normal modes of guest molecule and phonon for the quantized intermolecular normal modes of host matrix.
@@
with their order reflecting the hierarchy of energy scales associated with these states, as indicated in the bottom panel of Fig. 1.
@@
where $j$ is the integer index of the spin-orbital state and $S(T)$ denotes its spin multiplicity (singlet or triplet). In the following discussion of electronic and vibronic transitions, we will first omit the nuclear spin states, as they primarily manifest through hyperfine interactions with electron spins and are relevant only when probing the triplet manifold, such as in optically detected magnetic resonance experiments. We will revisit these states in Section II.3.
Figure 2: (a) The energy levels of a single molecule embedded in a matrix. The state |$\ket{S_{0(1)};\bm{\nu}_{v},\bm{\nu}_{p}}$ belongs to the ground (excited) singlet state manifold and has vibrational occupations in the intramolecular |$\ket{\bm{\nu}_{v}}=\Pi_{i}\ket{\nu_{v}^{(i)}}$ and intermolecular vibrational modes |$\ket{\bm{\nu}_{p}}=\Pi_{i}\ket{\nu_{p}^{(i)}}$. The indices $\nu_{v(p)}^{(i)}=0,1\ldots$ denote the vibrational quantum number of the $i^{\text{th}}$ vibron (phonon) mode. Similarly, |$\ket{T_{1};\bm{\nu}_{v},\bm{\nu}_{p}}$ represents the states in the triplet state manifold. (b) PAH molecules commonly used in quantum optics experiments and their range of approximated zero-phonon-line wavelengths [45].
@@ II.1.1 Ground and excited states
The ground state of the system |$\ket{S_{0};\bm{\nu}_{v}=\mathbf{0},\bm{\nu}_{p}=\mathbf{0}}$ is a singlet state, in which all $\pi$-orbitals are filled and all electron spins are paired. The excitation of one electron to a $\pi^{*}$ orbital leaves the molecule in the lowest singlet excited state |$\ket{S_{1};\bm{\nu}_{v}=\mathbf{0},\bm{\nu}_{p}=\mathbf{0}}$. Alternatively, if this excitation is accompanied by a spin-flip, the molecule reaches the lowest triplet excited state |$\ket{T_{1};\bm{\nu}_{v}=\mathbf{0},\bm{\nu}_{p}=\mathbf{0}}$. The energy of |$\ket{T_{1};\bm{\nu}_{v}=\mathbf{0},\bm{\nu}_{p}=\mathbf{0}}$ is typically lower than that of |$\ket{S_{1};\bm{\nu}_{v}=\mathbf{0},\bm{\nu}_{p}=\mathbf{0}}$ by the exchange energy. The transition energies of these two excited states to the ground state lie in the optical frequency domain and range from ultraviolet (UV) to near-infrared (NIR) wavelengths, depending on the size and geometry of the molecule. The triplet state |$\ket{T_{1};\bm{\nu}_{v}=\mathbf{0},\bm{\nu}_{p}=\mathbf{0}}$ consists of three sublevels, resulting from the anisotropic magnetic dipole-dipole interactions of two unpaired electrons, which remain non-degenerate even in the absence of an external magnetic field. This is called zero-field splitting[46]. For PAH molecules emitting in the UV to visible range, the zero-field splitting can reach a few gigahertz [47]. Moreover, coupling of the electron spin to nuclear spins in the molecule and the matrix leads to hyperfine splittings, which lie in the radio-frequency range.
@@
In addition, each manifold includes vibron states |$\ket{S(T),\bm{\nu}_{v}\neq\mathbf{0},\bm{\nu}_{p}=\mathbf{0}}$ corresponding to librational and bond vibrations of the guest molecule, with frequencies that can reach up to $100$ THz. The coupling between electronic and intramolecular vibrational states (electron-vibron coupling) gives rise to the well-known Franck-Condon physics, where the displacement of the excited state potential energy surface along a normal-mode coordinate enables transitions between vibronic states in different manifolds. The transition probabilities, i.e., Franck-Condon factors, are related to the Huang-Ryhs factor of each normal mode [48]. For certain PAH molecules, the simple Franck-Condon picture can be further complicated by the Herzberg-Teller effect, where couplings between the electronic and nuclear degrees of freedom are not parametric [49, 50]. Furthermore, non-adiabatic couplings between the electronic and nuclear degrees of freedom result in non-radiative decay from the excited state manifold via electron-vibron couplings [51, 52]. This could lead to non-unity quantum efficiencies of photon emission [53].
@@
Besides, phonon states |$\ket{S(T),\bm{\nu}_{v}=\mathbf{0},\bm{\nu}_{p}\neq\mathbf{0}}$ typically fall within the $10$ kHz - $5$ THz range and lead to the deformation of molecular orbitals. The resulting electron-phonon couplings yield various optical transitions between different manifolds involving vibrons and phonons, similar to the Franck-Condon physics discussed earlier. Depending on the frequency range, these vibrational states exhibit lifetimes in the range of milliseconds to picoseconds due to different vibrational relaxation mechanisms [54].
@@ II.1.2 Zero-phonon-line transition
In summary, a single molecule embedded in a solid-state host exhibits a rich energy structure supporting a multitude of well-defined transitions ranging from radio to optical frequencies. Next, we delve into details of these molecular degrees of freedom and discuss their quantum mechanical characters.
@@ II.2 Vibrational States
The optomechanical character of molecules was first discovered by Raman [59], as the optical excitation of molecules yields light scattering to new frequencies determined by the intramolecular vibrational energies. In the context of molecular quantum optics, this optomechanical character has traditionally been regarded as a source of decoherence due to fast vibrational relaxation processes. Yet, if vibrational states can be engineered to have long lifetimes, they could be utilized as a resource in quantum technologies.
Figure 3: (a) Phonon density of states of naphthalene and anthracene crystals (top), and intramolecular vibrations of a perylene molecule (bottom), shown by red, black, and green lines, respectively, plotted up to $12$ THz. For clarity, only $10$ out of $90$ vibrational modes of perylene that lie in this spectral range are shown. The ab initio calculations are done using Vienna Ab initio Simulation Package [60, 61, 62, 63] with Perdew-Burke-Ernzerhof functional [64] and Tkatchenko-Scheffler dispersion correction [65]. We used the Phonopy python package [66] to calculate the phonon density of states for a $2\times 4\times 2$ supercell. (b) Schematic of acoustic and optical phonons in an anthracene crystal. The crystal axes are denoted by $a$ and $b$. (c) Illustrations of local and pseudolocal modes in the host-guest system. (d) Vibrational relaxation mechanisms for an intramolecular vibration of a guest molecule in a host environment. The symbols $\omega_{\nu_{v}}$ and $\omega_{\text{max}}$ denote the intramolecular vibrational frequency of the state |$\ket{S_{0};[0,\ldots,0,1,0\dots,0],\bm{\nu}_{p}=\bm{0}}$ and maximal phonon frequency in the host matrix, respectively.